# Runtime Stealthy Perception Attacks against DNN-based Adaptive Cruise Control Systems

### v2

@@ 6.4 Evading Existing Defense Methods
We assess four widely used open-source defense methods that only rely on model input transformation without the need for re-training, including adding Gaussian noise [69], JPEG compression [70], reducing image color bit-depth [71], and using spatial median smoothing [71]. We evaluate the attack success rates in causing hazards under each defense method with various parameter settings while also considering the effect of input transformations on the benign or attack-free image frames and maintaining the baseline safety of the ACC system.
@@
As shown in Fig. 15, the experimental results reveal that JPEG compression and bit-depth reduction methods effectively reduce the attack success rate in causing hazards under specific parameter configurations. However, these methods fall short in maintaining the ACC’s safety by leading the benign image frames to cause hazards. In instances where the benign cases do not lead to hazards, the attack hazard rate is at 100%. On the other hand, the incorporation of Gaussian noise or median smoothing reduces the ACC’s lead vehicle detection accuracy. These methods are ineffective in mitigating CA-Opt attacks (hazard rate stays at 100% for all configurations), while also causing hazards for benign image frames.
Sensor Fusion. An alternative defense against adversarial patches could involve a fusion of an independent sensor (e.g., Lidar or radar) with the camera data for LVD predictions. While Lidar is reported prohibitively expensive for Level-2 AV [11], the fusion of radar and camera data fails to prevent controller misbehavior or collisions in our experiments (see Appendix A). This might be due to the common Kalman filter implementation of sensor fusion, which assumes the measurement noise follows a zero-mean Gaussian distribution. However, such design is vulnerable to perturbations with magnitudes less than one standard deviation of Gaussian noise (see Eq. 7) as reported in [18].
7 Discussion
Sim-to-Real Gap. Addressing the sim-to-real gap in AV security literature is challenging due to the risks and costs of real-road tests. In this paper, we aim to narrow this gap by developing a realistic experimental platform that integrates production ADAS control software, a physical-world simulator, and well-designed safety interventions with high-risk driving scenarios designed based on the NHTSA report [57]. Moreover, we evaluate the sim-to-real transfer possibility using an actual vehicle and a publicly available dataset. However, limitations exist, such as the fixed model and thresholds used for the human driver simulator that may impact evaluation results. Also, additional firmware and safety checks might be deployed in real cars, which can further limit the effectiveness of the proposed attack.
@@ 7 Discussion
8 Related Work
Adversarial Attacks on DNN. Many works have explored the vulnerability of DNN against adversarial attacks by adding adversarial physical/digital patches or stickers [9, 10, 72, 21, 73, 12, 11, 13, 74]. However, most of these works focus on altering the prediction class or probability or lane line position, which do not apply to attacks against ACC. Moreover, they rely on off-line optimization of attack value, neglecting the impact of attack timing. In contrast, our work introduces a novel runtime perception attack method against DNN-based ACC systems, employing a combined knowledge and data-driven approach that considers both attack timing and value for enhanced effectiveness. The only work [15] studied ACC vulnerability focuses on the physical attack without considering dynamic changes at runtime attack, which is also not scalable to many vehicles.
Security Analysis of AVs. Great efforts have also been made in studying the security of AVs, such as the security of Lidar [75], GPS [76], radar [77], camera [78], lane detection [11, 13], multiple objects tracking [18, 29, 28], control software [17], and safety mechanisms [16]. To the best of our knowledge, this paper is the first analysis of the security of Level-2 production ACC systems under stealthy safety-critical attack by considering three levels of safety interventions by constraint checking, human driver, and AEB/FCW and addressing unique challenges (Section 3.2).
9 Conclusion
This paper proposes a novel runtime stealthy attack strategy against DNN-based ACC systems, consisting of (i) a control-theoretic method for finding the most critical system contexts for launching the attacks to maximize the chance of safety hazards and (ii) an optimization-based image perturbation method for efficient generation and injection of adversarial patches to the DNN input to cause ACC control software misbehavior and hazards as soon as possible before being detected or mitigated by the ADAS safety mechanisms or human driver. Experimental results on an enhanced closed-loop simulation platform, a publicly available driving dataset, and an actual vehicle demonstrate the effectiveness of our proposed approaches for the strategic selection of attack start times and durations and the generation of the stealthy adversarial patches for perturbing DNN-based lead vehicle detection in a production Level-2 ADAS. This study also provides insights into the development of future ADAS that are robust against safety-critical attacks and the importance of interventions by the drivers and basic safety mechanisms for preventing attacks.
@@ Appendix A Sensor Fusion
Fig. 16 (Top) shows an example of predictions of the relative distance to the lead vehicle from the fusion of the camera and radar measurements. We see that the radar and camera predictions agree well most of the time. Also, the error between the fusion predictions and the ground truth relative distance (based on positions of vehicles in the simulator) becomes smaller as the Ego vehicle approaches the lead vehicle (an RMSE of 0.81m after 3000 control cycles in the figure). Sensor fusion also helps reduce the errors in fusion predictions under attacks as shown in Fig. 16 (Bottom) and discussed in Section 6.4, even though it fails to prevent collisions in the end.
@@ Appendix B Malware Installation
Fig. 17 shows an example set of steps taken for establishing remote access and downloading a malicious code repository on a vehicle running OpenPilot, the open-source production ACC system from Comma.ai [8].
Figure 17: An example of malware installation.
@@ Appendix D Stealthiness User Study
Our study included 30 participants who were asked to sit on the driver’s side of an autonomous vehicle, parked in a parking lot, equipped with OpenPilot ADAS (see Section 2.3). Each participant went through different trials of pre-recorded videos displayed on the ADAS monitor and answered a series of questions after each trial using a Qualtrics survey [82]. All participants had driving experience and 40% of them had autonomous driving experience.
@@
Stealthiness. We create five video sets by introducing adversarial patches into a pre-recorded highway scenario using CA-Random, CA-FGSM, and CA-Opt methods with three stealthiness levels ($\lambda=10^{-4}$, $\lambda=10^{-3}$, $\lambda=10^{-2}$, as detailed in Section 5.3.1). We present these videos on the ADAS monitor and ask participants whether they notice any abnormal scenarios that prompt them to assume control of the vehicle to avoid potential risk or danger. For a detailed examination, we extract an image frame from each video at the same frame index and zoom in to reveal more intricate details, followed by posing identical questions to the participants.
Results of the user study are illustrated in Fig. 19. It is evident that patches generated by the CA-Random attack are conspicuous to the majority of participants (>75%), whether observed in images or videos. In comparison, patches generated by the CA-FGSM exhibit lower visibility than those produced by the CA-Random attacks. In CA-Opt attacks, the takeover rate diminishes with a rise in stealthiness level or $\lambda$ value. Specifically, when $\lambda$ is set at $10^{-2}$ and $10^{-3}$, the takeover rates are below 20% for patch images and are 0% for patch videos. These findings suggest that adversarial patches at $\lambda=10^{-2}$ and $\lambda=10^{-3}$ exhibit nearly imperceptible characteristics to human drivers, particularly in image frames when not zoomed in.
@@ Appendix E Evaluation of Fake Video Attacks
Fake Video Attack. We rerun the simulations for each scenario introduced in Section 5.1 and replace the real-time camera frames with the selected fake video when the Ego vehicle approaches a similar position indicated by the fake video. Experimental results show that this attack causes hazards in 100% simulations and in 72.6% of simulations the Ego vehicle drives to the neighbor lane without any collisions. This is because lane lines in the fake video do not exactly overlap with the ones in the actual video.
@@
Fake Video Attack with Safety Interventions. To further evaluate the performance of fake video attacks with safety mechanisms, we rerun the experiments by launching the proposed driver intervention 2.5 seconds (average reaction time) after the attack. Experimental results show that all the attacks are successfully prevented. Therefore, we do not further test the fake video attack while enabling AEBS and constraint checking (see Section 4).
5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v3

@@ VI-D Evading Existing Defense Methods
We assess four widely used open-source defense methods that only rely on model input transformation without the need for re-training, including adding Gaussian noise [94], JPEG compression [33], reducing image color bit-depth [93], and using spatial median smoothing [93]. We evaluate the attack success rates in causing hazards under each defense method with various parameter settings while considering the effect of input transformations on the benign or attack-free image frames to maintain the baseline safety of the ACC system.
@@
As shown in Fig. 17, JPEG compression and bit-depth reduction methods effectively reduce the attack success rate in causing hazards under specific parameter configurations. However, these methods fall short in maintaining the ACC’s safety by leading the benign image frames to cause hazards. In instances where the benign cases do not lead to hazards, the attack hazard rate is at 100%. On the other hand, the incorporation of Gaussian noise or median smoothing reduces the ACC’s LV detection accuracy. These methods are ineffective in mitigating CA-Opt attacks (hazard rate stays at 100% for all configurations), while also causing hazards for benign frames.
Sensor Fusion. An alternative defense against adversarial patches could involve integrating independent sensors like Lidar or radar with camera data for LVD predictions. However, Lidar is too costly for Level-2 AVs [76], and our tests found that radar-camera fusion did not prevent ACC misbehavior or collisions (see Appendix A). This may be because of the use of Kalman filters in sensor fusion, which assumes measurement noise is zero-mean Gaussian and are vulnerable to perturbations smaller than one standard deviation of this noise [44]. In addition, sensor fusion outputs a weighted summation of radar and camera predictions, which cannot completely eliminate deviations caused by erroneous camera predictions, particularly when they significantly deviate from the ground truth. In some production ACC, camera predictions typically carry more weight. The vulnerability of sensor fusion was also reported in previous works [40]
VII Discussion
Sim-to-Real Gap. Addressing the sim-to-real gap in AV security literature is challenging due to the risks and costs of real-road tests. In this paper, we tried to narrow this gap by developing a realistic experimental platform that integrates production ADAS control software, a physical-world simulator, and well-designed safety interventions with high-risk driving scenarios designed based on the NHTSA report [67]. Moreover, we evaluate the sim-to-real transfer possibility using an actual vehicle, a model lead car, and a publicly available dataset. However, limitations exist, such as the fixed model and thresholds used for the human driver simulator that may impact evaluation results. Also, additional firmware and safety checks might be deployed in real cars, which can further limit the effectiveness of the proposed attack.
@@ VII Discussion
VIII Related Work
Adversarial Attacks on DNN. Many works have explored the vulnerability of DNN against adversarial attacks by adding adversarial physical/digital patches or stickers [55, 51, 41, 37, 65, 36, 76, 84, 90]. However, most of these works focus on altering the prediction class or probability or lane line position, which do not apply to attacks against ACC. Moreover, they rely on off-line optimization of attack value, neglecting the impact of attack timing. In contrast, our work introduces a novel runtime perception attack method against DNN-based ACC systems, employing a combined knowledge and data-driven approach that considers both attack timing and value for enhanced effectiveness. The only other work on ACC [39] focused on the physical attacks without considering dynamic changes at runtime, which is not scalable to many vehicles.
Security Analysis of AVs. Great efforts have also been made in studying the security of AVs, such as the security of Lidar [75], GPS [81], radar [47], camera [74], lane detection [76, 84], multiple objects tracking [44, 45, 58], control software [97], and safety mechanisms [60]. To the best of our knowledge, this paper is the first analysis of the security of Level-2 production ACC systems under stealthy safety-critical attack by considering three levels of safety interventions by constraint checking, human driver, and AEB/FCW and addressing unique challenges (Section III-B).
IX Conclusion
This paper proposes a novel runtime stealthy attack strategy against DNN-based ACC systems, consisting of (i) a control-theoretic method for finding the most critical system contexts for launching the attacks to maximize the chance of safety hazards and (ii) an optimization-based image perturbation method for efficient generation and injection of adversarial patches to the DNN input to cause ACC control misbehavior and hazards as soon as possible before being detected or mitigated by the ADAS safety mechanisms or human driver. Experiments on a production Level-2 ADAS using an enhanced closed-loop simulation platform, a publicly available driving dataset, and an actual vehicle demonstrate the effectiveness of our approach in improving attack success rate and stealthiness compared to different baselines. This study also provides insights into the development of future ADAS that are robust against safety-critical attacks and the importance of interventions by the drivers and basic safety mechanisms for preventing attacks.
@@ Appendix A Sensor Fusion
Fig. 18 (Top) shows an example of predictions of the relative distance to the lead vehicle from the fusion of the camera and radar measurements. We see that the radar and camera predictions agree well most of the time. Also, the error between the fusion predictions and the ground truth relative distance (based on positions of vehicles in the simulator) becomes smaller as the Ego vehicle approaches the lead vehicle (an RMSE of 0.81m after 3000 control cycles in the figure). Sensor fusion also helps reduce the errors in fusion predictions under attacks as shown in Fig. 18 (Bottom) and discussed in Section VI-D, even though it fails to prevent collisions in the end.
@@ Appendix B Malware Installation
Fig. 19 shows an example set of steps taken for establishing remote access and downloading a malicious code repository on a vehicle running OpenPilot, the open-source production ACC system from Comma.ai [29]. To change live camera image frames at runtime, we use a technique called monkey patching in Python, which is used for dynamically modifying or extending the behavior of an existing code at runtime or hooking a function without changing the source code[42]. For example, as shown in Fig. 19, the send() system library that is called by the camera call back function for sending the received camera frames to the DNN model can be wrapped by a malicious version adv_send() that implements the attack.
Figure 19: An example of malware installation (Top) and malicious code execution using monkey patching technique (Bottom).
@@ Appendix D Stealthiness User Study
Our study included 30 participants who were asked to sit on the driver’s side of an autonomous vehicle, parked in a parking lot, equipped with OpenPilot ADAS (see Section II-C). Each participant went through different trials of pre-recorded videos displayed on the ADAS monitor and answered a series of questions after each trial using a Qualtrics survey [8]. All participants had driving experience and 40% of them had autonomous driving experience.
@@
Stealthiness. We create five video sets by introducing adversarial patches into a pre-recorded highway scenario using CA-Random, CA-APGD, and CA-Opt methods with three stealthiness levels ($\lambda=10^{-4}$, $\lambda=10^{-3}$, $\lambda=10^{-2}$, as detailed in Section V-C1). We present these videos on the ADAS monitor and ask participants whether they notice any abnormal scenarios that prompt them to assume control of the vehicle to avoid potential risk or danger. For a detailed examination, we extract an image frame from each video at the same frame index and zoom in to reveal more intricate details, followed by posing identical questions to the participants.
Results of the user study are illustrated in Fig. 21. It is evident that patches generated by the CA-Random attack are conspicuous to the majority of participants (>75%), whether observed in images or videos. In comparison, patches generated by the CA-APGD exhibit lower visibility than those produced by the CA-Random attacks. In CA-Opt attacks, the takeover rate diminishes with a rise in stealthiness level or $\lambda$ value. Specifically, when $\lambda$ is set at $10^{-2}$ and $10^{-3}$, the takeover rates are below 20% for patch images and are 0% for patch videos. These findings suggest that adversarial patches at $\lambda=10^{-2}$ and $\lambda=10^{-3}$ exhibit nearly imperceptible characteristics to human drivers, particularly in image frames when not zoomed in.
@@ Appendix E Evaluation of Fake Video Attacks
Fake Video Attack. We rerun the simulations for each scenario introduced in Section V-A and replace the real-time camera frames with the selected fake video when the Ego vehicle approaches a similar position indicated by the fake video. Experimental results show that this attack causes hazards in 100% simulations and in 72.6% of simulations the Ego vehicle drives to the neighbor lane without any collisions. This is because lane lines in the fake video do not exactly overlap with the ones in the actual video.
@@
Fake Video Attack with Safety Interventions. To further evaluate the performance of fake video attacks with safety mechanisms, we rerun the experiments by launching the proposed driver intervention 2.5 seconds (average reaction time) after the attack. Experimental results show that all the attacks are successfully prevented. Therefore, we do not further test the fake video attack while enabling AEBS and constraint checking (see Section IV).
Appendix F Testing on a Real-world Dataset.
We perform an evaluation using the comma2k19 dataset [77], a publicly available dataset with over 33 hours of California’s 280 highway commute. The dataset comprises 2019 segments, each lasting one minute, covering a 20km highway section, collected using OpenPilot hardware.
Table XI compares the CA-Opt attack with CA-Random for different distances between the Ego and lead vehicles. CA-Random has an average deviation of 0.15m, which does not significantly impact ACC system outputs or cause hazards as the ACC system typically keeps a following distance larger than 4 meters in the absence of attacks. In contrast, when the Ego vehicle is close to the lead vehicle (less than 20m), CA-Opt achieves the highest deviation of 18.65m, showcasing the effectiveness of the proposed objective function (Section III-C2) in generating optimal perturbations with substantial impact.
TABLE XI: Performance of CA-Opt attack vs. CA-Random in deviating DNN-based lead vehicle position predictions using comma2k19 dataset.
Attack Metric Longitudinal Deviation in DNN Prediction (m)
0-20 20-40 40-60 60-80 80+ All
CA-Random Avg. 0.70 0.41 0.28 0.99 0.08 0.15
Std 0.69 1.03 1.44 1.93 2.67 2.56
CA-Opt Avg. 18.65 16.15 14.52 8.65 3.73 4.91
Std 6.96 4.83 4.95 2.82 3.03 3.35